OpenStudy (anonymous):

whats the added # to make a perfect square: k^2+11k

OpenStudy (cwrw238):

121/4: (k + 11/2)(k + 11/2) = k^2 + 11k + 121/4

OpenStudy (anonymous):

@cwrw238 can you break it down for me?

OpenStudy (cwrw238):

(k + 11/2)(k + 11/2 = k^2 + 11/2* k + 11/2 * k + 11/2 * 11/2 = k^2 + 11k + 121/4

OpenStudy (cwrw238):

you need to divide the 11k by 2 and square the result

OpenStudy (cwrw238):

- * sorry : square the 11/2

OpenStudy (anonymous):

\[11/2^{2} is 2.75 or 11/4 ??\]

OpenStudy (cwrw238):

no its (11/2)^2 = 11^2 / 2^2 = 121/4

OpenStudy (anonymous):

so the # the make a perfect square is 121/4 ??